OpenStudy (anonymous):

Help please! I'd really appreciate it. :) Earl purchased a living room set for $3,592 using a 12-month deferred payment plan. The interest rate after the introductory period is 21.80%. A down payment of $275 is required as well as a minimum monthly payment of $112. What is the balance after the introductory period if only the minimum payment is made until then? $4,004.92 $2,839.23 $4,116.92 $2,772.92

OpenStudy (anonymous):

@jim_thompson5910 Hi there! Do you think you could possibly help me with this? :)

jimthompson5910 (jim_thompson5910):

what do you have so far?

OpenStudy (anonymous):

Well, honestly I'm stuck on what formula I should use.

jimthompson5910 (jim_thompson5910):

how much money is loaned to Earl

OpenStudy (anonymous):

$3,592

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

Do I have to add the down payment to that also?

jimthompson5910 (jim_thompson5910):

he makes a down payment

jimthompson5910 (jim_thompson5910):

that reduces the amount that is loaned to him

jimthompson5910 (jim_thompson5910):

example: if the item is $100 and he makes a $10 down payment, then he has to pay off 100-10 = 90 dollars. So 90 dollars is loaned to him

OpenStudy (anonymous):

$3,317 then?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

Now do I start to utilize the monthly payments he paid throughout the introductory period (which adds up to $1,344)?

jimthompson5910 (jim_thompson5910):

it says "The interest rate after the introductory period is 21.80%" notice the keyword "after"

jimthompson5910 (jim_thompson5910):

so I'm going to assume that the interest rate before the introductory period is 0%

OpenStudy (anonymous):

Right -- so we technically don't even have to use the interest rate for this problem?

jimthompson5910 (jim_thompson5910):

which essentially means that if he can pay off the $3,317 in its entirety before the interest rate goes up, then he won't pay a penny of interest

jimthompson5910 (jim_thompson5910):

yeah so I'm thinking we simply do 3317 - 1344 = 1,973 that's the balance after he's made those 12 payments of $112

OpenStudy (anonymous):

Hmm . . . what's the next step then? It seems like that's the number it's asking for, but it's not one of the choices.

jimthompson5910 (jim_thompson5910):

yeah somehow interest will be added onto that (since he didn't get to $0 before the introductory period is up)

jimthompson5910 (jim_thompson5910):

ok I figured it out

OpenStudy (anonymous):

Ok, great!

jimthompson5910 (jim_thompson5910):

so if Earl pays off the entire balance before the 12 months are up, then he doesn't pay a penny of interest

jimthompson5910 (jim_thompson5910):

however, if he has a nonzero balance by the time the 12 months are up, he's charged interest on the original balance (unfair I know, but that's how it works) so because 3317-1344 = 1973 is nonzero, he still has a balance which accrues interest. This means that we take the 3317 and use it as principle and have the interest compound A = P(1+r/n)^(n*t) A = 3317 (1+0.2180/12)^(12*1) A = 4,116.91564539528 A = 4,116.92 So when those 12 months are up, the balance is $4,116.92 If he doesn't make any payments, then the balance sticks to $4,116.92 But he made 12 payments of $112, so we subtract off 12*112 = 1344 to get 4,116.92 - 1,344 = 2,772.92

jimthompson5910 (jim_thompson5910):

so you can see why it's a good idea for Earl to pay off all the balance by month 12

OpenStudy (anonymous):

Wow! That's actually pretty unfair, it totally negates the whole "0% interest" for the introductory period in this case due to his nonzero amount . . . I feel bad for Earl lol :P Anyway, thank you so much for breaking it down! I appreciate your assistance to the fullest. :D

jimthompson5910 (jim_thompson5910):

yeah it's deferred in a sense that it's saying "we'll put off the interest rate for 12 months IF you can pay off the balance. If not, then we'll go back to day 1 and recompute interest just as if the 0% interest rate never happened"

jimthompson5910 (jim_thompson5910):

so on the surface, it looks like a good deal but only if you can pay it all off by 12 months